Write the equation of the line which has a slope of eight ninths
8
9 and passes through ​(9,2​). Write the answer in​ slope-intercept form.

Answer by macston(5194)

m=slope; x=9; y=2
y=mx+b
y=(8/9)x+b
2=(8/9)9+b
2=8+b
-6=b
.
Final equation:
y=(8/9)x-6